Good day,
What is the half-life of a first-order reaction with a rate constant of 4.30×10−4 s−1?
1
Expert's answer
2020-06-04T10:38:47-0400

For the first-order chemical reaction A"\rightarrow" products, the rate law is:

"[A] = [A]_0e^{-kt}" ,

where "k" is the rate constant and "t" is the time.

The half-life of a chemical reaction is the time that it takes to decrease the concentration of the reagent A by a factor of two:

"\frac{[A]}{[A]_0} = 0.5 = e^{-kt_{1/2}}" .

Therefore, the half-life of the reaction is:

"t_{1/2} = \frac{\text{ln2}}{k} = \frac{\text{ln2}}{4.30·10^{-4}} = 1612" s.

Answer: 1612 s.